OpenStudy (anonymous):

I have no idea how to solve this: Determine the quadrant in which an arc of s=20 on the unit circle would terminate.

OpenStudy (loser66):

the first one

OpenStudy (loser66):

I mean the first quadrant

OpenStudy (anonymous):

How did you get that?

OpenStudy (loser66):

I guess, hihihihi...

OpenStudy (loser66):

unit circle has radius 1, so that the perimeter is 2pi = 2*3.1415 = 6.283 so that each quadrant has the length 6.283/4 = 1.57075 and 20 = 3*6.283 + 1.151 That means I go around the circle 3 times and then go a 1.151 arc length. And this amount is at the first quadrant. that's what I thought. Not sure whether it's valid or not. @jim_thompson5910

OpenStudy (anonymous):

thanks!!